OpenStudy (anonymous):

Express in lowest terms. 27a^3-1/9a^2-6a+1 @p

OpenStudy (campbell_st):

lol... to funny....

OpenStudy (anonymous):

@ChmE

OpenStudy (anonymous):

@Jonask

OpenStudy (anonymous):

@sue101

OpenStudy (anonymous):

@brookester6

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

factorise both nominator and denomnator\[27a^3-1=(3a)^3-1=(3a-1)(9a^2+3a+1)\]

OpenStudy (anonymous):

ummm...what grade r u in??

OpenStudy (anonymous):

lowest terms and simplest term do mean the same right

OpenStudy (anonymous):

\[\frac{ (3a-1)(9a^2+3a+1) }{ (3a-1)(3a-1) }=\frac{ 9^2+3a+1 }{ 3a-1 }\]